You want to buy a new sports coupe for $78,500, and the finance office at the dealership has quoted you a loan with an AR of 4.9 percent for 60 months to buy the car. What will your monthly payments be? What is the effective rate on this loan?
Answer #1

Information provided:

Present value= $78,500

Time= 60 months

Interest rate= 4.9%/12= 0.4083% per month

The monthly payment is calculated by entering the below in a financial calculator:

PV= -78,500

N= 60

I/Y= 0.4083

Press the CPT key and PMT to compute the amount of monthly payment.

The value obtained is 1,463.0560.

Therefore, the amount of monthly payment is $1,463.06.

Effective annual rate is calculated using the below formula:

EAR= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

EAR= (1+0.049/12)^12-1

        = 1.050116-1

        = 0.050116*100= 5.0116%

Therefore, the effective annual rate is 5.01%.

In case of any query, kindly comment on the solution